Home‎ > ‎Crew 142‎ > ‎

1107-GreenHabReport

GreenHab Report
11/07/2014
Carmel Johnston

Current Temperature:
75
Minimum Temperature:
45
Maximum Temperature:
93
Current Humidity:
15
Minimum Humidity:
13
Maximum Humidity:
20
Hydro Tank Level:

Hydro Tank P H:

Hydro Tank Temp F:

Hydro Tank Over Flow:

Plant Report:
2014-2015 season lettuce: starts were transported from the Hab to
GreenHab in the morning. No
sprouting has occurred yet. They were again transported back to the
Hab this evening.

Crew 142 Martian soil project: Seeds are beginning to germinate,
mostly visable by seeds pushing
themselves up through the soil. Seeds will be covered up with their
own soil in the morning to
ensure proper root formation.
Watering Activity Notes:
2014-2015 season lettuce: starts were moistened lightly with a spray
bottle to ensure not
overwatering.

Crew 142 Martian soil project: All crew 142 experiment trays were
deemed to be sufficiently
moist, so watering was postponed. Watering will continue tomorrow
morning at 5mL (radishes) and
0.5mL (chard) once per day unless it is deemed that there adequate moisture.
Heater Activity Notes:
Heater was turned on at 9pm and turned off at 8:30 am. Heating resumed
at 6:30pm.
Fan Activity Notes:
Fan ran from 8:30 am until 5pm.
Hydroponic Time:

Hydroponic Duration:

Hydroponic Kw H:

Led Lamp Time:

Led Lamp Duration:

Led Lamp Kw H:

Maintenance Report:
An empty bucket has been borrowed from the GreenHab to allow for a
backup toilet if we have
troubles with the septic again tonight. The bucket will be double
lined with garbage bags to
ensure no contamination of human waste.
Supplies Request: